OpenStudy (anonymous):

Jenny bought a coat at a "20% off" sale and saved $35. What was the original price of the coat?

OpenStudy (anonymous):

is/of = %/100

OpenStudy (anonymous):

35/x = 20/100

OpenStudy (anonymous):

Cross multiply: 3500 = 20x x = $175

OpenStudy (anonymous):

let the original price = $ x 20% off = (20/100)*x = 20x/100 ATQ, 20x ---- = 35 100 35 * 100 x = ---------- = 35*5 = 175 20 so original price was $175